MinMax Corp has the following capital structure: 55% equity (giving a return of 9%), 10% preferred shares (with a yield of 6%), and 35% debt (with a coupon rate of 10% and yield to maturity of 6.5%). If there are no taxes, what is the firm’s WACC?

Testbank, Question 11 In the same question above, what would be the WACC if taxes are included at a corporate tax rate of 40%? 10.333% 9.050% 7.825% 6.915%
Answer #1
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=6.5*0.35+9*0.55+6*0.1
WACC =7.83%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 6.5*(1-0.4)
= 3.9
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.9*0.35+9*0.55+6*0.1
WACC =6.915%